# Real exponential sums over primes and prime gaps

### v2

@@ 2. The adaptation of Newman’s proof
If we wanted to continue following very closely in Newman’s footsteps, we should now use some kind of algebraic identity (of the type described in Step $8$) to finish the proof of the formulas 1.1 and 1.2. But this would lead us to a problem of commuting an infinite double series (explained in section 4) that I honestly don’t know how to work around. Even so, we can complete the proof of the formulas 1.1 and 1.2. How?
@@
For $s\in\mathbb{C}$ such that $\Re(s)>0$, define the functions
Obviously these functions are analytic on $\Re(s)>0$.
Proposition 2.14.
The function $\Psi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$ if, and only if, the function $\Xi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$.
Proof.
$(\Rightarrow)$ By assumption, there is an analytic function $s\mapsto F(s)$, defined on an open set containing $\Re(s)\geq 0$, such that
$\Psi(s)-\displaystyle\frac{1}{s}=F(s),$
for all $s\in\mathbb{C}$ such that $\Re(s)>0$, that is
$\Xi(s)-\displaystyle\frac{1}{s}=(1-\lambda)F((1-\lambda)s),$
$(\Leftarrow)$ By assumption, there is an analytic function $s\mapsto G(s)$, defined on an open set containing $\Re(s)\geq 0$, such that
$\Xi(s)-\displaystyle\frac{1}{s}=G(s),$
for all $s\in\mathbb{C}$ such that $\Re(s)>0$, that is,
for all $s\in\mathbb{C}$ such that $\Re(s)>0$. This shows that $\Psi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$. ∎
Step 9
In turn, the same happens with the function $\tau(s)$.
Proposition 2.15.
The function $\Xi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$ if, and only if, the function $\tau(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$.
Proof.
Note that for all $s\in\mathbb{C}$ such that $\Re(s)>0$ holds
$\begin{split}\Xi(s)-\displaystyle\frac{1}{s}&=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{\lambda}\exp[s(p^{1-\lambda}-1)]}-\displaystyle\frac{1}{s}\\ &=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{\lambda}\exp(sp^{1-\lambda})\exp(-s)}-\displaystyle\frac{1}{s}\\ &=\exp(s)\tau(s)-\displaystyle\frac{1}{s}\\ &=\exp(s)\left[\tau(s)-\displaystyle\frac{1}{s}-\left(\displaystyle\frac{1}{s\exp(s)}-\displaystyle\frac{1}{s}\right)\right].\end{split}$
Step 10
This is our last step, and the point where all our work culminates. First, for $s\in\mathbb{C}$ such that $\Re(s)>0$, define the function
$T(s)=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{2-\lambda}\exp(sp^{1-\lambda})}.$
Obviously this function is analytic on $\Re(s)>0$. Now, consider the following well-known theorem.
@@
initially defined on $\Re(s)>0$, extends to a meromorphic function on $\Re(s)>-\displaystyle\frac{1}{2}$ that is analytic on $\Re(s)\geq 0$.
@@
Let $\Omega(s)$ be the extension referred in the theorem 2.16. The main idea to finish the proof of formulas 1.1 and 1.2 is to calculate the Mellin transform of the function $T(s)$, rewrite the obtained expression in a convenient way, and then calculate the inverse Mellin transform. This will reveal to us the analytic continuation of the function $\tau(s)-\displaystyle\frac{1}{s}$.
@@
The function $\tau(s)-\displaystyle\frac{1}{s}$, initially defined for $\Re(s)>0$, extends analytically to $\Re(s)>-1$.
@@
As mentioned above, calculating the Mellin transform of the function $T(s)$ for $\Re(s)>0$, we will obtain
$\begin{split}\mathscr{M}(z)&=\mathscr{M}\left(\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{2-\lambda}\exp(sp^{1-\lambda})}\right)\\ &=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{2-\lambda}}\mathscr{M}[\exp(-sp^{1-\lambda})]\\ &=\displaystyle\sum_{p}\displaystyle\frac{(1-\lambda)\log(p)}{p^{2-\lambda}}\cdot\displaystyle\frac{\Gamma(z)}{p^{(1-\lambda)z}}\\ &=(1-\lambda)\Gamma(z)\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{1+(1-\lambda)(z+1)}}\\ &=(1-\lambda)\Gamma(z)\Phi(1+(1-\lambda)(z+1)).\end{split}$
Note that this function is meromorphic on $\Re(z)>-1$ with a simple pole at $z=0$, not because of the function $\Phi(1+(1-\lambda)(z+1))$, which is analytic on $\Re(z)>-1$, but because of the function $\Gamma(z)$ with its simple pole at $z=0$. Using the information from theorem 2.16, we can write
@@
$\begin{split}\mathscr{M}(z)&=(1-\lambda)\Gamma(z)\left[\displaystyle\frac{1}{(1-\lambda)(z+1)}+\Omega((1-\lambda)(z+1))\right]\\ &=\displaystyle\frac{\Gamma(z)}{z+1}+(1-\lambda)\Gamma(z)\Omega((1-\lambda)(z+1)).\end{split}$
Now notice that this function has exponential decay along vetical lines, because $\Gamma(z)$ does, and that the function $\Omega((1-\lambda)(z+1))$ doesn’t change that (by lemma 3.8). Therefore, we can get back the function $T(s)$ by computing the inverse Mellin transform of $\mathscr{M}(z)$, and write
@@
$\begin{split}T(s)&=\mathscr{M}^{-1}(\mathscr{M}(z))\\ &=\mathscr{M}^{-1}\left(\displaystyle\frac{\Gamma(z)}{z+1}+(1-\lambda)\Gamma(z)\Omega((1-\lambda)(z+1))\right)\\ &=\mathscr{M}^{-1}\left(\displaystyle\frac{\Gamma(z)}{z+1}\right)+(1-\lambda)\mathscr{M}^{-1}[\Gamma(z)\Omega((1-\lambda)(z+1))]\\ &=\exp(-s)-s\cdot\Gamma(0,s)+(1-\lambda)\underbrace{\mathscr{M}^{-1}[\Gamma(z)\Omega((1-\lambda)(z+1))]}_{\text{analytic on }\Re(s)>-1},\end{split}$
for all $s\in\mathbb{C}$ such that $\Re(s)>0$. Differentiating twice, one gets
@@
$\begin{split}\tau(s)&=T^{\prime\prime}(s)\\ &=\displaystyle\frac{d^{2}}{ds^{2}}\left(\exp(-s)-s\cdot\Gamma(0,s)\right)+(1-\lambda)\displaystyle\frac{d^{2}}{ds^{2}}\left(\mathscr{M}^{-1}[\Gamma(z)\Omega((1-\lambda)(z+1))]\right)\\ &=\displaystyle\frac{\exp(-s)}{s}+(1-\lambda)\underbrace{\displaystyle\frac{d^{2}}{ds^{2}}\left(\mathscr{M}^{-1}[\Gamma(z)\Omega((1-\lambda)(z+1))]\right)}_{\text{analytic on }\Re(s)>-1}.\end{split}$
Finally, subtracting $\frac{1}{s}$ on either side of this equation, we get
$\tau(s)-\displaystyle\frac{1}{s}=\displaystyle\frac{\exp(-s)-1}{s}+(1-\lambda)\underbrace{\displaystyle\frac{d^{2}}{ds^{2}}\left(\mathscr{M}^{-1}[\Gamma(z)\Omega((1-\lambda)(z+1))]\right)}_{\text{analytic on }\Re(s)>-1}.$
But since the function $s\mapsto\displaystyle\frac{\exp(-s)-1}{s}$ extends analytically to $\mathbb{C}$, we are done. ∎
@@
Proof.
Since $\tau(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)>-1$, then $\Psi(s)-\displaystyle\frac{1}{s}$ extends analytically to $\Re(s)\geq 0$, thus proving the validity of formulas 1.1 and 1.2. ∎

### v3

@@ 2. The adaptation of Newman’s proof
If we wanted to closely follow in Newman’s footsteps, we would now have to use some kind of algebraic identity (of the type described in step $8$) to finish the proof of the formulas 1.1 and 1.2. But, honestly, I don’t know how to do that. Even so, we can complete the proof of formulas 1.1 and 1.2. How?
@@
Consider the following well-known theorem.
@@
initially defined on $\Re(s)>0$, extends to a meromorphic function on $\Re(s)>-1/2$ that is analytic on $\Re(s)\geq 0$.
@@
Let $\Omega(s)$ be the extension referred in the theorem 2.14. The main idea to finish the proof of formulas 1.1 and 1.2 is to expand the function
$\lambda\mapsto\displaystyle\frac{p^{s+1-\lambda}}{\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}$
in its Taylor series at $\lambda=1$, and inserting the information from theorem 2.14 into the expression obtained. For now, consider the following definition.
Definition 2.15.
Let $n\in\mathbb{N}$. We say that a polynomial $f(x)\in\mathbb{Q}[x]$ is special (or $n$-special, when we want to indicate its degree) if $\deg(f)=n$ and if
$\mathscr{F}[f(x),s]=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+1}f(s\log(p))}{p^{s+1}},$
Let $n\in\mathbb{N}$ and $f(x)=a_{0}+a_{1}x+\cdots a_{n}x^{n}\in\mathbb{Q}[x]$. Of course $\mathscr{F}[f(x),s]$ is defined for $\Re(s)>0$, since
$\begin{split}\mathscr{F}[f(x),s]&=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+1}[a_{0}+a_{1}s\log(p)+\cdots a_{n}s^{n}\log(p)^{n}]}{p^{s+1}}\\ &=\displaystyle\sum_{j=0}^{n}a_{j}s^{j}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+j+1}}{p^{s+1}},\end{split}$
but this does not necessarily imply that $\mathscr{F}[f(x),s]$ extends analytically to $\Re(s)\geq 0$. Let’s look at two examples.
Example 2.16.
$f(x)=x$ is not special, because
$\mathscr{F}[x,s]=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{2}s\log(p)}{p^{s+1}}=s\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{3}}{p^{s+1}}=s\Phi^{\prime\prime}(s+1)=s\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime\prime}=s\Omega^{\prime\prime}(s)+\displaystyle\frac{2}{s^{2}}$
has a pole of order $2$ at $s=0$.
Example 2.17.
f(x)=x2f(x)=x-2 is $1$-special, because
$\begin{split}\mathscr{F}[x-2,s]&=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{2}[s\log(p)-2]}{p^{s+1}}\\ &=s\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{3}}{p^{s+1}}-2\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{2}}{p^{s+1}}\\ &=s\Phi^{\prime\prime}(s+1)+2\Phi^{\prime}(s+1)\\ &=s\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime\prime}+2\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime}\\ &=s\Omega^{\prime\prime}(s)+\displaystyle\frac{2}{s^{2}}+2\Omega^{\prime}(s)-\displaystyle\frac{2}{s^{2}}\\ &=s\Omega^{\prime\prime}(s)+2\Omega^{\prime}(s)\end{split}$
is analytic on $\Re(s)\geq 0$.
@@
The polynomials
  • $3x^{2}-16x+12$
  • $x^{3}-10x^{2}+22x-8$
  • $15x^{4}-240x^{3}+1040x^{2}-1248x+240$
  • $3x^{5}-70x^{4}+500x^{3}-1248x^{2}+912x-96$
  • $63x^{6}-2016x^{5}+21420x^{4}-90944x^{3}+146664x^{2}-69120x+4032$
  • $9x^{7}-378x^{6}+5586x^{5}-36008x^{4}+102088x^{3}-113136x^{2}+35568x-1152$
are all specials.
@@
Let’s just do the first case; the rest is left to the reader. Indeed,
@@
$\begin{split}\mathscr{F}[3x^{2}-16x+12,s]&=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{3}[3s^{2}\log(p)^{2}-16s\log(p)+12]}{p^{s+1}}\\ &=3s^{2}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{5}}{p^{s+1}}-16s\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{4}}{p^{s+1}}+12\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{3}}{p^{s+1}}\\ &=3s^{2}\Phi^{\prime\prime\prime\prime}(s+1)+16s\Phi^{\prime\prime\prime}(s+1)+12\Phi^{\prime\prime}(s+1)\\ &=3s^{2}\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime\prime\prime\prime}+16s\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime\prime\prime}+12\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{\prime\prime}\\ &=3s^{2}\Omega^{\prime\prime\prime\prime}(s)+\displaystyle\frac{72}{s^{3}}+16s\Omega^{\prime\prime\prime}(s)-\displaystyle\frac{96}{s^{3}}+12\Omega^{\prime\prime}(s)+\displaystyle\frac{24}{s^{3}}\\ &=3s^{2}\Omega^{\prime\prime\prime\prime}(s)+16s\Omega^{\prime\prime\prime}(s)+12\Omega^{\prime\prime}(s).\end{split}$
is analytic on $\Re(s)\geq 0$. ∎
Proposition 2.19.
Let $f(x)\in\mathbb{Q}[x]$ such that $\deg(f)=n\geq 1$, and write
$f(x)=a_{0}+a_{1}x+\cdots+a_{n}x^{n}.$
Then $f(x)$ is special if, and only if,
$\sum_{k=0}^{n}\alpha_{k}\cdot(n+k)!=0.$
Proof.
Simply calculate $\mathscr{F}[f(x),s]$ and note that
@@
$\begin{split}\mathscr{F}[f(x),s]&=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+1}f(s\log(p))}{p^{s+1}}\\ &=\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+1}}{p^{s+1}}\displaystyle\sum_{k=0}^{n}\alpha_{k}s^{k}\log(p)^{k}\\ &=\displaystyle\sum_{k=0}^{n}\alpha_{k}s^{k}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{n+k+1}}{p^{s+1}}\\ &=\displaystyle\sum_{k=0}^{n}\alpha_{k}s^{k}(-1)^{n+k}\Phi^{(n+k)}(s+1)\\ &=\displaystyle\sum_{k=0}^{n}\alpha_{k}s^{k}(-1)^{n+k}\left[\Omega(s)+\displaystyle\frac{1}{s}\right]^{(n+k)}\\ &=\displaystyle\sum_{k=0}^{n}\alpha_{k}s^{k}(-1)^{n+k}\left[\Omega^{(n+k)}(s)+\displaystyle\frac{(-1)^{n+k}\cdot(n+k)!}{s^{n+k+1}}\right]\\ &=\displaystyle\sum_{k=0}^{n}(-1)^{n+k}\alpha_{k}s^{k}\Omega^{(n+k)}(s+1)+\displaystyle\frac{1}{s^{n+1}}\displaystyle\sum_{k=0}^{n}\alpha_{k}\cdot(n+k)!\end{split}$
is analytic on $\Re(s)\geq 0$ if, and only if, $\sum_{k=0}^{n}\alpha_{k}\cdot(n+k)!=0$. ∎
Now, consider the sequence of polynomials $\{f_{j}(x)\}_{j\geq 0}$ in $\mathbb{Q}[x]$ defined recursively by
(*) $f_{0}(x)=1\qquad\qquad\text{and}\qquad\qquad f_{j+1}(x)=f_{j}(x)-x\displaystyle\sum_{k=0}^{j}\displaystyle\binom{j}{k}\displaystyle\frac{f_{j-k}(x)}{k+2},$
for all $j\geq 0$. It is easy to see that these polynomials have the following properties:
  • $\deg(f_{j})=j$, for all $j\geq 0$;
  • $f_{j}(0)=1$, for all $j\geq 0$;
  • the leading coefficient of $f_{j}(x)$ is $\left(-\displaystyle\frac{1}{2}\right)^{j}$, for all $j\geq 0$.
These polynomials will play a fundamental role in the next (and final) step of proving formulas 1.1 and 1.2. They appear as Taylor’s expansion of the expression
$\frac{p^{s+1-\lambda}}{\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}=\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(1-\lambda)^{j}\log(p)^{j}f_{j}(s\log(p))}{j!}$
at $\lambda=1$, through the use of the exponential formula (see [18]).
If we write each polynomial $f_{j}(x)$ in the form
$f_{j}(x)=\displaystyle\sum_{k=0}^{j}\alpha_{j,k}\cdot x^{k},$
this same formula allows us to calculate the coefficients recursively, using the following expressions:
$\alpha_{j,0}=1,\qquad\forall\ j\geq 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\alpha_{j,j}=\left(-\displaystyle\frac{1}{2}\right)^{j},\qquad\forall\ j\geq 0,$
and
$\alpha_{j+1,k+1}=\displaystyle\frac{(j+1)(k+2)\alpha_{j,k+1}-(j+1)\alpha_{j,k}}{j+k+2},\qquad\forall\ j\geq 1,\ \forall\ k\in\{0,\ldots,j-1\}.$
Now, we need the following result.
Proposition 2.20.
Consider the sequence of polynomials $\{f_{j}(x)\}_{j\geq 0}$ defined above. Then $f_{j}(x)$ is special, for all $j\geq 1$.
Proof.
We will prove this by induction on $j$, using the criterion established in proposition 2.19. If $j=1$, then
$\alpha_{1,0}\cdot 1!+\alpha_{1,1}\cdot 2!=1\cdot 1!+\left(-\displaystyle\frac{1}{2}\right)\cdot 2!=1-1=0.$
Now, suppose that
$\sum_{k=0}^{j}\alpha_{j,k}\cdot(j+k)!=0.$
We need to show that
$\sum_{k=0}^{j+1}\alpha_{j+1,k}\cdot(j+1+k)!=0.$
Indeed, $\sum_{k=0}^{j+1}\alpha_{j+1,k}\cdot(j+1+k)!=$
@@
$\begin{split}{}&=\alpha_{j+1,0}\cdot(j+1)!+\displaystyle\sum_{k=1}^{j}\alpha_{j+1,k}\cdot(j+1+k)!+\alpha_{j+1,j+1}\cdot(2j+2)!\\ &=1\cdot(j+1)!+\displaystyle\sum_{k=0}^{j-1}\alpha_{j+1,k+1}\cdot(j+k+2)!+\left(-\displaystyle\frac{1}{2}\right)^{j+1}\cdot(2j+2)!\\ &=(j+1)!+\displaystyle\sum_{k=0}^{j-1}\displaystyle\frac{(j+1)(k+2)\alpha_{j,k+1}-(j+1)\alpha_{j,k}}{j+k+2}\cdot(j+k+2)!+\left(-\displaystyle\frac{1}{2}\right)^{j+1}\cdot(2j+2)!\\ &=(j+1)!+(j+1)\displaystyle\sum_{k=0}^{j-1}(k+2)\alpha_{j,k+1}\cdot(j+k+1)!-(j+1)\displaystyle\sum_{k=0}^{j-1}\alpha_{j,k}\cdot(j+k+1)!\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\left(-\displaystyle\frac{1}{2}\right)^{j+1}\cdot(2j+2)!\\ &=(j+1)!+(j+1)\displaystyle\sum_{k=0}^{j-2}(k+2)\alpha_{j,k+1}\cdot(j+k+1)!+(j+1)\cdot(j+1)\alpha_{j,j}\cdot(2j)!\\ &\qquad\qquad\qquad\qquad-(j+1)\alpha_{j,0}\cdot(j+1)!-(j+1)\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k+1)!+\left(-\displaystyle\frac{1}{2}\right)^{j+1}\cdot(2j+2)!\\ &=(j+1)!+(j+1)\displaystyle\sum_{k=1}^{j-1}(k+1)\alpha_{j,k}\cdot(j+k)!+(j+1)^{2}\cdot\left(-\displaystyle\frac{1}{2}\right)^{j}\cdot(2j)!\\ &\qquad\qquad\qquad\qquad-(j+1)\cdot(j+1)!-(j+1)\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k+1)!+\left(-\displaystyle\frac{1}{2}\right)^{j+1}\cdot(2j+2)!\\ &=(j+1)!\cdot[1-(j+1)]+(j+1)\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k)!\cdot[(k+1)-(j+k+1)]\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\left(-\displaystyle\frac{1}{2}\right)^{j}\cdot(2j)!\cdot\left[(j+1)^{2}+\left(-\displaystyle\frac{1}{2}\cdot(2j+1)\cdot(2j+2)\right)\right]\\ &=-j\cdot(j+1)!-j(j+1)\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k)!+\left(-\displaystyle\frac{1}{2}\right)^{j}\cdot(2j)!\cdot[-j\cdot(j+1)]\\ &=-j\cdot(j+1)\cdot\left[j!+\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k)!+\left(-\displaystyle\frac{1}{2}\right)^{j}\cdot(2j)!\right]\\ &=-j\cdot(j+1)\cdot\left[\alpha_{j,0}\cdot j!+\displaystyle\sum_{k=1}^{j-1}\alpha_{j,k}\cdot(j+k)!+\alpha_{j,j}\cdot(2j)!\right]\\ &=-j\cdot(j+1)\cdot\displaystyle\sum_{k=0}^{j}\alpha_{j,k}\cdot(j+k)!=0,\end{split}$
by the induction hypothesis. ∎
Step 9
Proposition 2.21.
The function $\Psi(s)-\displaystyle\frac{1}{s}$, initially defined for $\Re(s)>0$, extends analytically to $\Re(s)\geq 0$.
Proof.
Following the strategy we discussed at the beginning of step $8$, we write
$\Psi(s)=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{\lambda}\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{s+1}}\cdot\displaystyle\frac{p^{s+1-\lambda}}{\exp\left[s\left(\displaystyle\frac{p^{1-\lambda}-1}{1-\lambda}\right)\right]}.$
By the exponential formula (see [18]), we have
$\Psi(s)=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{s+1}}\cdot\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(1-\lambda)^{j}\log(p)^{j}f_{j}(s\log(p))}{j!},$
where this sequence of polynomials $\{f_{j}(x)\}_{j\geq 0}$ is defined by (*2). Since this series converges locally uniformly in $\Re(s)>0$, we can change the order of the double sum and write
$\begin{split}\Psi(s)&=\displaystyle\sum_{j=0}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{s+1}}\\ &=\displaystyle\sum_{p}\displaystyle\frac{\log(p)}{p^{s+1}}+\displaystyle\sum_{j=1}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{p}\displaystyle\frac{\log(p)^{j+1}f_{j}(s\log(p))}{p^{s+1}}\\ &=\Phi(s+1)+\displaystyle\sum_{j=1}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\cdot\mathscr{F}[f_{j}(x),s].\end{split}$
Writing each $f_{j}(x)$ as $f_{j}(x)=\displaystyle\sum_{k=0}^{j}\alpha_{j,k}\cdot x^{k}$ as before, one has
$\begin{split}\Psi(s)&=\Phi(s+1)+\displaystyle\sum_{j=1}^{\infty}\displaystyle\frac{(1-\lambda)^{j}}{j!}\displaystyle\sum_{k=0}^{j}(-1)^{j+k}\alpha_{j,k}\cdot s^{k}\cdot\Omega^{(j+k)}(s)\\ &=\Phi(s+1)+\displaystyle\sum_{j=1}^{\infty}(\lambda-1)^{j}\cdot\Omega^{(j)}(s)\cdot\displaystyle\sum_{k=0}^{\lfloor j/2\rfloor}\displaystyle\frac{\alpha_{j-k,k}}{(j-k)!}\cdot\left(\displaystyle\frac{s}{1-\lambda}\right)^{k}.\end{split}$
Finally, subtracting $1/s$ on either side of this equation, we get
$\begin{split}\Psi(s)-\displaystyle\frac{1}{s}&=\Phi(s+1)-\displaystyle\frac{1}{s}+\displaystyle\sum_{j=1}^{\infty}(\lambda-1)^{j}\cdot\Omega^{(j)}(s)\cdot\displaystyle\sum_{k=0}^{\lfloor j/2\rfloor}\displaystyle\frac{\alpha_{j-k,k}}{(j-k)!}\cdot\left(\displaystyle\frac{s}{1-\lambda}\right)^{k}\\ &=\Omega(s)+\displaystyle\sum_{j=1}^{\infty}(\lambda-1)^{j}\cdot\Omega^{(j)}(s)\cdot\displaystyle\sum_{k=0}^{\lfloor j/2\rfloor}\displaystyle\frac{\alpha_{j-k,k}}{(j-k)!}\cdot\left(\displaystyle\frac{s}{1-\lambda}\right)^{k}\\ &=\displaystyle\sum_{j=0}^{\infty}(\lambda-1)^{j}\cdot\Omega^{(j)}(s)\cdot\displaystyle\sum_{k=0}^{\lfloor j/2\rfloor}\displaystyle\frac{\alpha_{j-k,k}}{(j-k)!}\cdot\left(\displaystyle\frac{s}{1-\lambda}\right)^{k}.\qed\end{split}$